An old dilapidated bungalow has just been bought and the new owner wanted to give a
fresh look by constructing a rectangular garden on its backyard. The length of a
rectangular garden in 5 m more than its width and the area is 50 m2
. Find the length,
width and perimeter of the garden.

Answer by ikleyn(52786)

.

If w meters is the width, the the length is (w+5) meters.


The area equation is


    w * (w+5) = 50.    (*)


You can solve this quadratic equation


    w^2 + 5w - 50 = 0


using factoring


    (w-5)*(w+10) = 0,


which gives you a unique positive solution w= 5.


Or you can guess the answer in 5 seconds, looking into equation (*).


The width is 5 meters, the length is  5+5 = 10 meters, and the perimeter is

 
      5 + 10 + 5 + 10 = 30 meters.

Solved.